# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

### v7

@@ 4.3.3 Results
Tables 25 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset. As expected, combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy. However, min-entropy ($H_{\infty}$) remains stubbornly low. For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ (Shannon) but saturates at only 21 bits of $H_{\infty}$. Interestingly, we find that omitting certain correlated sensors sometimes does not reduce min-entropy at all. For the Relay dataset in Table 3, the combination (Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the full set’s $8.092$ bits. Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$. This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
@@
Figure 3: Entropy values versus the number of joint sensor modalities.
[…]
The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality. Combining signals increases the apparent capacity for unpredictability, but correlations between sensors means that the min-entropy from a large sensor ensemble is not be substantially higher than a reduced subset thereof. We note that standards implore the use of min-entropy as a safer, worst-case metric [26, 33]. It measures how close the distribution is to collapsing around the single most probable outcome that an adversary will target first. To the best of our knowledge, the entropy collapse from highly correlated sensor modalities has not been elicited in existing work. We provide the full sensor combination results in our open-source repository.2
@@
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 48.061 25.089 17.116 11.008
(Acc.{x,y,z}, Gyro.{y,z}) 39.403 22.835 15.999 10.113
(Acc.{x,y,z}, Gyro.{x,y}) 39.886 21.456 15.332 9.900
(Acc.{x,y,z}, Gyro.{x,z}) 40.222 21.470 15.040 9.411
(Acc.{y,z}, Gyro.{x,y,z}) 40.228 21.306 14.698 9.274
(Acc.{x,z}, Gyro.{x,y,z}) 40.293 21.260 14.575 9.154
(Acc.{x,y,z}, Gyro.y) 31.228 18.804 13.893 9.065
(Acc.{x,y}, Gyro.{x,y,z}) 40.273 21.155 14.208 8.721
(Acc.{x,y,z}, Gyro.z) 31.564 18.775 13.833 8.682
(Acc.{y,z}, Gyro.{y,z}) 31.570 18.583 13.400 8.386
@@
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 45.794 25.036 15.725 8.092
(Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 44.795 24.963 15.334 7.859
(Acc., Grav., Light, Lin. Acc., Mag., Rot. Vec.) 38.385 21.590 14.783 7.702
(Acc., Grav., Gyro., Light, Mag., Rot. Vec.) 37.026 21.001 14.553 7.702
(Grav., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 36.092 20.794 14.450 7.673
(Acc., Light, Lin. Acc., Mag., Rot. Vec.) 37.385 21.517 14.428 7.469
(Acc., Gyro., Light, Mag., Rot. Vec.) 36.026 20.924 14.270 7.465
(Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 35.092 20.720 14.147 7.439
(Acc., Grav., Gyro., Light, Lin. Acc., Mag.) 41.094 22.794 14.437 7.338
(Acc., Grav., Light, Mag., Rot. Vec.) 29.617 17.555 13.244 7.312
@@
Modalities $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 83.662 36.998 33.682 23.926
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 76.907 34.258 31.510 23.835
(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.246 31.509 23.835
(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.820 34.243 31.508 23.829
(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.229 31.501 22.945
(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 71.384 32.586 30.190 22.945
(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 68.568 34.218 31.500 22.945
(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 68.650 34.215 31.499 22.945
(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 67.214 32.575 30.189 22.945
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 76.618 33.816 31.035 21.915
@@
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 158.601 82.301 39.320 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 148.995 78.624 39.276 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 148.759 78.477 39.272 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 148.587 78.380 39.249 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 148.952 78.135 39.235 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 139.153 74.800 39.175 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 138.981 74.703 39.128 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 138.745 74.556 39.117 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 139.346 74.458 39.100 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 139.109 74.311 39.087 21.289
@@ 5 Evaluation
This section critically evaluates our sensor entropy findings using existing literature and relevant standards. We then discuss potential mitigations such as randomness extractors, outline inherent limitations of our study and sensor-based approaches, and propose avenues for future enhancements.
@@ 5.1 Discussion
Prior studies have claimed or implied that mobile sensors are suitable data sources for security-critical applications. Much work has relied on model evaluation metrics as a proxy for evaluating such claims [13, 14, 4, 2], with a smaller subset using more established entropy metrics [25, 24]. However, analyses have demonstrated that individual sensors confer very little entropy, especially under worst-case assumptions [17, 16, 15, 7]. Our results, derived from an analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.
@@
Single sensors yield low entropy. Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits). Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g. some single-sensor readings yield only 1–3 bits of min-entropy). This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not fully capture.
Multiple sensors improve worst-case entropy, but not dramatically. While combining more modalities typically raises the upper bound (e.g. Hartley or Shannon entropy), our experiments reveal that min-entropy gains are far smaller than one might hope. In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution. Our findings place stricter bounds on the benefits and risks of mobile-sensor-based approaches. We estimate that even the most complicated multi-modal combination provide relatively little worst-case entropy. It is important to note that our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further. Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone source. They might be useful for tasks where strong assurances are not required, e.g. simple proximity checks; however, relying on sensor data as robust entropy sources is fraught with security risks.
What can be done? One might hope that cryptographic extractors (e.g. Von Neumann extractors or more advanced schemes [47, 48, 49]) could make low-entropy sensors suitable for usage. Extractors are designed to reduce bias in a noisy or skewed source; however, they cannot increase the total amount of unpredictability beyond the source’s intrinsic min-entropy. If combined sensor data provides, say, 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more. This means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance). Extraction enhances the quality of the original source if it has enough min-entropy, but it cannot elevate a source that fundamentally lacks it.
Another avenue is to introduce additional unpredictability through user interaction or deliberate environmental perturbation. For example, requiring the user to perform a random shaking gesture during a pairing protocol could inject extra entropy into the sensor readings. Prior studies have found that deliberate device movements can, indeed, increase the usable entropy of motion sensors, albeit only by a modest amount (on the order of a few bits) [15, 17]. While this boost is non-negligible, it remains far below the dozens or hundreds of bits typically desired for security-critical applications.
One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudo-random number generator (CSPRNG). By doing so, a hybrid design can mitigate bias in the raw sensor data and produce outputs that pass statistical randomness tests (see the designs by Suciu et al. [27] and Wallace et al. [29]). However, this approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings. For example, in matching similar motion traces or environmental conditions of two neighbouring devices [42, 13, 2, 3, 4, 1]. The desired property here is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices. Using CSPRNG-derived signals would make it difficult to distinguish whether signals originated from co-located devices using similarity-based metrics. This exposes a fundamental security paradox in existing work. We have shown that sensors have insufficient entropy, even when using multiple ones simultaneously. Yet, methods to increase that entropy would compromise the usefulness of schemes that rely on the similarity of one or more sensor signals.
We see no straightforward remedy for the entropy shortfall of mobile sensors. Extractors, user-assisted randomness, and entropy pools address diHoweverfferent problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution. Nevertheless, they do not compensate for fundamentally weak signals, which we have shown to be the case across a range of modalities. Sensor data and its derivatives may still be appropriate for low-security purposes. However, relying on it alone for security applications is highly inadvisable due to the serious risk of adversaries who exploit statistical biases to reduce the effective search space.
5.2 Limitations

### v8

@@ 4.3.3 Results
Tables 47 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset. As expected, combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy. However, min-entropy ($H_{\infty}$) remains stubbornly low. For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ but saturates at only 21 bits of $H_{\infty}$. Interestingly, we find that omitting certain correlated sensors sometimes does not reduce min-entropy at all. For the Relay dataset in Table 5, the combination (Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the full set’s $8.092$ bits. Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$. This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
@@
Figure 3: Average joint entropy vs. number of combined sensors modalities.
We plot the number of combined modalities for each dataset in Figure 3, showing the average (mean) joint entropy for different entropy measures. The data illustrates that, while $H$ values are tightly coupled initially, they diverge significantly as the number of joint modalities increases. All datasets demonstrated a trend of worst-case $H_{\infty}$ being significantly lower than other entropy values, often barely increasing relative to best-case metrics. For SHL and PerilZIS, for example, $H_{\infty}$ increases by approximately only 1–2 bits per modality, with $H_{0}$$H_{2}$ increasing far more significantly ($\Delta H_{0}\approx$ 10 bits and $\Delta H_{1}\approx$ 5 bits for every added modality). In Tables 47, we give the differences of $H_{1}$ and $H_{\infty}$ for the top combinations as $H_{1}-H_{\infty}$, quantifying the sharp decreases between Shannon and min-entropy as the number of combined modalities increases. We also use # to denote the number of modalities used in each row.
The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality. Combining signals increases the apparent capacity for unpredictability, but redundancies between sensors means that the min-entropy from a large sensor ensemble is not substantially higher than a reduced subset thereof. This aligns with the emphasis of using min-entropy for worst-case guarantees in modern standards [25, 24]. To our knowledge, the extent of entropy collapse under correlated mobile modalities has not been elicited at this scale. Full combination results are available in our repository (see Footnote 3).
@@
Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 6 48.061 25.089 17.116 11.008 14.081
(Acc.{x,y,z}, Gyro.{y,z}) 5 39.403 22.835 15.999 10.113 12.722
(Acc.{x,y,z}, Gyro.{x,y}) 5 39.886 21.456 15.332 9.900 11.556
(Acc.{x,y,z}, Gyro.{x,z}) 5 40.222 21.470 15.040 9.411 12.059
(Acc.{y,z}, Gyro.{x,y,z}) 5 40.228 21.306 14.698 9.274 12.032
(Acc.{x,z}, Gyro.{x,y,z}) 5 40.293 21.260 14.575 9.154 12.106
(Acc.{x,y,z}, Gyro.y) 4 31.228 18.804 13.893 9.065 9.739
(Acc.{x,y}, Gyro.{x,y,z}) 5 40.273 21.155 14.208 8.721 12.434
(Acc.{x,y,z}, Gyro.z) 4 31.564 18.775 13.833 8.682 10.093
(Acc.{y,z}, Gyro.{y,z}) 4 31.570 18.583 13.400 8.386 10.197
@@
Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 7 45.794 25.036 15.725 8.092 16.944
(Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 6 44.795 24.963 15.334 7.859 17.104
(Acc., Grav., Light, Lin. Acc., Mag., Rot. Vec.) 6 38.385 21.590 14.783 7.702 13.888
(Acc., Grav., Gyro., Light, Mag., Rot. Vec.) 6 37.026 21.001 14.553 7.702 13.299
(Grav., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 6 36.092 20.794 14.450 7.673 13.121
(Acc., Light, Lin. Acc., Mag., Rot. Vec.) 5 37.385 21.517 14.428 7.469 14.048
(Acc., Gyro., Light, Mag., Rot. Vec.) 5 36.026 20.924 14.270 7.465 13.459
(Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 5 35.092 20.720 14.147 7.439 13.281
(Acc., Grav., Gyro., Light, Lin. Acc., Mag.) 6 41.094 22.794 14.437 7.338 15.456
(Acc., Grav., Light, Mag., Rot. Vec.) 5 29.617 17.555 13.244 7.312 10.243
@@
Modalities # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 13 83.662 36.998 33.682 23.926 13.072
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 11 76.907 34.258 31.510 23.835 10.423
(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.246 31.509 23.835 10.411
(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.820 34.243 31.508 23.829 10.414
(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.229 31.501 22.945 11.284
(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 10 71.384 32.586 30.190 22.945 9.641
(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.568 34.218 31.500 22.945 11.273
(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.650 34.215 31.499 22.945 11.270
(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 9 67.214 32.575 30.189 22.945 9.630
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 11 76.618 33.816 31.035 21.915 11.901
@@
Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 22 158.601 82.301 39.320 21.289 61.012
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 148.995 78.624 39.276 21.289 57.335
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 148.759 78.477 39.272 21.289 57.188
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 21 148.587 78.380 39.249 21.289 57.091
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 20 148.952 78.135 39.235 21.289 56.846
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 19 139.153 74.800 39.175 21.289 53.511
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 138.981 74.703 39.128 21.289 53.414
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 138.745 74.556 39.117 21.289 53.267
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 19 139.346 74.458 39.100 21.289 53.169
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 19 139.109 74.311 39.087 21.289 53.022
4.3.4 Binning Sensitivity Analysis
As we discussed in §4.1 and §4.3.1, bin width is a key consideration in estimating entropy accurately and efficiently. Too few bins can lead to oversmoothing, where distinct features of the data distribution are lost, artificially deflating the entropy estimate. Conversely, too many bins can lead to sparsity and overfitting, where the analysis models the specific sample’s noise rather than the source’s true underlying distribution, inflating entropy. We performed a sensitivity analysis on the best-performing sensor combination (“All sensors”) for each of our four datasets. We re-calculated the joint entropy while varying the number of bins for each modality, starting from 5 bins and increasing up to 2048. Beyond that, the computational cost of processing the large state space became prohibitive. The results of this analysis are presented in Figure 4.
The graphs show a consistent trend across all datasets. For a low number of bins, all entropy measures are severely underestimated as expected. As the number of bins increases, the entropy rises sharply before tapering off and converging towards a stable plateau that approximates a logarithmic shape. (An exception is the min-entropy value for SHL, which converges after approximately 100 bins). Crucially, FD-selected bin counts fall within the stable region, supporting FD as a balanced choice between fidelity and tractability.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 4: Bin sensitivity analysis showing how entropy values vary by average bin width for the best-performing sensor combinations for each data set. Freedman–Diaconis (FD) values are also shown.
@@ 5 Evaluation
This section interprets our entropy findings for system design: we discuss the key takeaways, relate min-entropy to concrete guessing cost, discuss mitigations, and give recommendations for future research in the area.
@@ 5.1 Discussion
Prior studies have claimed that mobile sensors are suitable data sources for supporting security-critical applications. Much work has relied on model evaluation metrics as a proxy for evaluating such claims [7, 6, 21, 9], with a smaller subset using more established entropy metrics [36, 15]. Separate entropy analyses of sensors have demonstrated that individual sensors actually confer very little entropy, particularly under worst-case assumptions [28, 27, 26, 17]. Our results both challenge and extend these claims.
@@
Single sensors yield low entropy. Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits). Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g. some single-sensor readings yield only 1–3 bits of min-entropy). This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not capture at this scale.
Combining multiple sensors offers modest gains to worst-case min-entropy. While using multimodal sensors typically raises best- and average-case entropies ($H_{0}$, $H_{1}$), our experiments reveal that gains in worst-case min-entropy ($H_{\infty}$) are far smaller than one might expect. In many cases, $H_{\infty}$ plateaus, reflecting the dominance of a handful of high-probability joint outcomes. Even the most complex multimodal combinations in our datasets yield only modest worst-case entropy, well below the levels required by modern security standards. This also aligns with prior work showing that Shannon entropy can significantly overstate security [54, 55, 24, 25]. Hence, while mobile sensors can augment other authentication or key-generation processes, they do not suffice by themselves. Using multiple sensors may still improve usability and model accuracy, e.g. for coarse co-location checks, but they should not be treated as secure, unpredictable sources.
5.2 Translating Entropy Bounds Into Guessing Attacks
We sketch how low-entropy sensor data translates into concrete attack cost. Let $Z$ be the discretised joint sensor vector used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy. Ordering guesses by decreasing probability, the success probability after $q$ guesses satisfies
$\Pr[\text{success within }q]\;\leq\;q\cdot 2^{-H_{\infty}(Z)}.$ (10)
The expected number of guesses is $\approx 2^{H_{\infty}(Z)-1}$ [55].
Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching [7, 6, 9, 21, 35, 20, 19]. Let us consider an abstract authentication function, $\mathrm{Auth}(z)\to\{0,1\}$, which accepts or rejects an input vector $z\in Z$. An adversary’s task becomes easier through a similarity-based scheme: they only need to generate a candidate $z^{\prime}$ that is ‘close enough’ to a high-probability vector $z$. This means the effective success probability for each guess is, in fact, higher than the bound in Eq. 10 suggests. The best-performing combination in our study yielded $H_{\infty}\approx 24$ bits, corresponding to a search space of $2^{24}$ ($\approx$16.7 million) possibilities. An adversary may pre-compute the most likely vectors and submit them to $\mathrm{Auth}$, requiring only $2^{23}$ guesses on average. This is well within the capabilities of a moderately resourced adversary—taking seconds to milliseconds on a modern PC to enumerate—and is orders of magnitude weaker than the $2^{128}$ expected of standard cryptographic keys [23]. Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach [42]. In the worst cases, such as the single-sensor readings where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring approximately $2^{7}=128$ guesses on average. While rate-limiting the authentication function can mitigate this risk, it does not solve the root problem. If each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$. Rate limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
5.3 What can be done?
Our results highlight a core tension in sensor-based security design. Proposals for pairing, co-location inference, or shared-secret agreement typically depend on the similarity of sensor signals across devices. Yet the very regularities that enable similarity seemingly stem from low-entropy structure. In effect, these systems trade cryptographic unpredictability for application-level utility. The open question is whether a scheme can ever achieve both meaningful entropy and the cross-device similarity it depends on. We close by offering several recommendations for researchers in this area.
Evaluate the threat of adversarial samples
A persistent feature in the literature is the lack of rigorous analysis of min-entropy (and induced guesswork) under realistic acceptance regions and attacker priors [49, 7, 9, 10, 21, 8, 5, 6]. We urge the community to consider the resilience of proposals to adversarial samples, particularly when using ‘fuzzy’ matching methods such as supervised learning models, and to translate those estimates to concrete costs, such as expected guesses.
Quantify effects of deliberate interaction and induced phenomena
User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits [26, 28]. For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors. A key gap is to quantify how much $\Delta H_{\infty}$ such interventions can actually deliver across modalities. Our expectation is that any gains remain modest, but this trade-off between added entropy and practical feasibility remains an under-explored area of research.
Decouple sensor similarity from unpredictability
Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs. The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see [38, 40]). Cryptographic extractors (e.g. Von Neumann and later constructions [56, 57, 58]) or decorrelation techniques (e.g. PCA) may be considered to convert low-quality sensor noise into higher-entropy sources. However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy. That is, if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more. Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost. Future work in the area should focus on how entropy can be meaningfully increased while retaining application utility.
5.4 Scope and Limitations